# MS-Mix: Unveiling the Power of Mixup for Multimodal Sentiment Analysis

### v1

@@ 4.4. Verification Performance of MS-Mix
+ Manifold Mix 78.56 65.21 37.86 78.72 0.442 65.09
+ MultiMix 78.24 65.21 38.11 78.41 0.447 64.99
+ $\mathcal{P}ow$Mix* 79.04 - - 78.51 0.437 -
+ $\mathcal{P}ow$Mix 78.64 65.49 37.85 78.90 0.444 65.22
+ MS-Mix (ours) 78.77 65.79 38.73 78.62 0.441 65.48
MISA 76.81 63.24 37.42 76.39 0.467 63.47
+ Manifold Mix 77.19 64.01 36.29 77.05 0.456 63.64
+ MultiMix 77.50 64.18 37.06 77.18 0.437 63.98
+ $\mathcal{P}ow$Mix* 77.35 - - 76.97 0.441 -
+ $\mathcal{P}ow$Mix 77.53 64.03 37.51 77.16 0.439 64.06
+ MS-Mix (ours) 78.54 64.16 37.83 78.18 0.431 64.68
ALMT 76.81 63.02 34.35 76.81 0.449 62.75
+ Manifold Mix 76.37 64.99 36.32 76.66 0.460 63.59
+ MultiMix 77.68 62.63 38.95 78.03 0.433 64.32
+ $\mathcal{P}ow$Mix* 78.91 - - 79.13 0.429 -
+ $\mathcal{P}ow$Mix 78.17 63.06 37.53 78.24 0.431 64.28
+ MS-Mix (ours) 78.34 64.46 37.68 78.21 0.426 64.67
GLoMo 79.09 66.31 37.75 79.20 0.427 65.59
+ Manifold Mix 78.82 66.42 38.02 79.06 0.426 65.58
+ MultiMix 78.89 66.38 38.40 79.27 0.424 65.74
+ $\mathcal{P}ow$Mix 79.03 65.43 38.53 79.16 0.434 65.53
+ MS-Mix (ours) 79.28 67.69 38.06 79.52 0.422 66.14
As shown in Tables 2 to 4, MS-Mix achieves top performance, ranking first or second in all metrics compared to baseline and other mixup-based methods. It is particularly noteworthy that MS-Mix consistently outperforms all competing approaches on the comprehensive average (Avg.) metric across every backbone model.
@@
The consistent outperformance of MS-Mix across diverse backbones and datasets stems from its core design principles, which directly address key limitations in existing mixup-based augmentation methods such as Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), and $\mathcal{P}ow$Mix (Georgiou et al., 2024). Unlike these methods, which often rely on random or offline mixing strategies, MS-Mix incorporates three interconnected mechanisms tailored for MSA: First, the SASS strategy prevents the blending of semantically contradictory samples (e.g., happy and angry) by filtering pairs based on emotional similarity in the latent space. In contrast, the comparative methods perform mixing between random sample pairs regardless of emotional coherence, often resulting in ambiguous or noisy samples that mislead training. By ensuring semantic and label consistency, SASS significantly enhances the quality of augmented data. Second, the SIG mixing module adaptively determines modality-specific mixing ratios based on emotional salience, rather than using random or fixed values as in Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), or leveraging modality importance without emotional context like $\mathcal{P}ow$Mix. This allows MS-Mix to dynamically weight the contribution of each modality, leading to more discriminative mixed features and improved robustness of learned cross-modal representations. Third, the SAL serves as a regularization term that aligns predicted sentiment distributions with ground-truth labels, enhancing prediction consistency across modalities and providing additional supervisory signal. This component is unique to MS-Mix and offers a critical advantage over comparative methods, which lack explicit constraints on emotion-level distribution alignment. As a result, SAL not only stabilizes training but also strengthens the model’s resistance to overfitting.
To further demonstrate the effectiveness of MS-Mix, as shown in Fig. 3, we employed t-SNE (Maaten and Hinton, 2008) to visualize the textual modal feature distributions of both original and mixed features on the MOSEI dataset, using the MISA model as the backbone. The t-SNE visualization demonstrates that features generated by MS-Mix exhibit significantly clearer decision boundaries and more distinct cluster separation compared to those produced by $\mathcal{P}$owMix. This improvement indicates that MS-Mix not only preserves the semantic integrity of the original feature space but also generates more discriminative intermediate representations.
Overall, MS-Mix effectively addresses the semantic confusion and label mismatch issues that hindered previous methods. By integrating three core strategies, our method generates discriminative features with corresponding labels, significantly enhancing the overall quality of augmented data.
4.5. Ablation and Hyperparameter Sensitivity Analysis
To evaluate the impact of key hyperparameters and individual algorithmic components of MS-Mix, we performed ablation studies on three core components using the LMF (Liu and Shen, 2018) model on the MOSI (Zadeh et al., 2016) and SIMS (Yu et al., 2020) datasets. Additionally, hyperparameter sensitivity analyses on the MOSEI dataset (Zadeh et al., 2018b), using the MISA (Hazarika et al., 2020) and GLoMo (Zhuang et al., 2024) models.
In our ablation study, we integrated Manifold Mixup (Verma et al., 2019) into the classic LMF (Liu and Shen, 2018) backbone as the baseline to evaluate the contribution of each proposed component: SASS, SIG, and SAL. Each module was selectively enabled/disabled during training to assess its individual impact on the performance of MS-Mix. Note that since the computation of SAL relies on the output of the SIG module, it cannot be used independently. Results presented in Table 5 show that each of the three components contributes noticeably to performance gains over the baseline, underscoring the importance of every module. Furthermore, by incrementally incorporating these components, we observed progressive improvements in performance, demonstrating not only the individual efficacy of each module but also the significance of synergistic work.
@@ 4.5. Ablation and Hyperparameter Sensitivity Analysis
Table 5. Results of the ablation experiments using the LMF method on the MOSI database. Bold: Best performance.
@@
82.16 80.90 82.12 80.81 42.42 36.88 0.893 67.55
@@
Fig. 4. The impact of different parameter values. (a) Accuracy under different similarity thresholds $\delta$. (b) Accuracy under different combinations of $\xi_{1}$ and $\xi_{2}$. (c) Accuracy under different $\alpha$. (d) Accuracy under different attention heads $h$. :
@@
To comprehensively analyze the sensitivity of key hyperparameters, we conducted extensive experiments using both the top-performing GLoMo model (Zhuang et al., 2024) and the classical MISA architecture (Hazarika et al., 2020) on the largest English dataset MOSEI (Zadeh et al., 2018b) and the Chinese benchmark SIMS (Yu et al., 2020). As illustrated in Fig. 4, our evaluation reveals how critical hyperparameters affect model performance. Using the MISA model on MOSEI, we first examined the similarity selection threshold $\delta$ (Fig. 4(a)). The results demonstrate that at $\delta=0.2$, the method maintains an optimal balance, incorporating sufficient samples while effectively filtering out emotionally contradictory ones, thus achieving peak performance. Notably, even under suboptimal $\delta$ configurations, MS-Mix consistently surpasses Manifold Mixup (Verma et al., 2019). We further evaluated the loss weights $\xi_{1}$ and $\xi_{2}$ with the GLoMo model on MOSEI (Fig. 4(b)). The experimental outcomes confirm the rationality of our weight selection, showing that $\xi_{1}=0.7$ and $\xi_{2}=0.5$ yield the best performance. Additionally, we assessed the shape parameter $\alpha$ of the Beta distribution using MISA on MOSEI (Fig. 4(c)), finding $\alpha=2.0$ to be optimal. The influence of the number of attention heads $h$ was also investigated across datasets (Fig. 4(d)): $h=4$ delivers the best results on MOSEI, while $h=6$ is most effective on SIMS. These findings collectively validate the rationale behind our parameter configurations.
4.6. Occlusion Experiment
@@
In the random occlusion experiment, we compared the performance of our method against baseline approaches (Verma et al., 2019; Venkataramanan et al., 2023; Georgiou et al., 2024) under both clean and noisy data conditions across MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018b), and SIMS (Yu et al., 2020). Specifically, we randomly masked out 0% to 40% of the input data during training of the LMF (Liu and Shen, 2018) model to investigate the effect of mixup-based augmentation on model robustness.
@@
Fig. 5. The performance of different mixup-based methods at different occlusion ratios on three datasets. :
@@
As illustrated in Fig. 5, the experimental results demonstrate that the introduction of noise influences model performance across all datasets. A noteworthy phenomenon emerges in which moderate levels of random masking (e.g., 10% or 20%) unexpectedly improve performance, rather than degrade it. Beyond these optimal masking ratios, however, model performance deteriorates. Despite these variations, MS-Mix consistently achieves the highest average accuracy across all evaluated conditions, demonstrating its advantages in optimizing the model for handling noisy data.
5. CONCLUSION
@@
In this paper, we propose MS-Mix, a novel and adaptive data augmentation framework specifically designed for MSA. Unlike traditional mixup-based methods, which often rely on random and offline mixing strategies, MS-Mix introduces three key emotion-aware mechanisms to enhance the semantic consistency of augmented samples. First, the SASS module ensures that only emotionally consistent samples are mixed, effectively avoiding mixing feature pairs with contradictory emotions. Second, the SIG mixing module adaptively determines mixing ratios of each modality based on emotional salience, promoting the generation of samples with clearer classification boundaries. Third, the SAL aligns the predicted emotion distributions with ground-truth labels through the KL-divergence minimization, serving as an effective regularizer to improve generalization. Empirical evaluation of three benchmark datasets and six diverse backbone architectures has shown that MS-Mix achieves state-of-the-art performance.
@@
In the future, we will consider extending MS-Mix to other multimodal tasks, such as mental health monitoring and human-computer interaction, which could leverage its ability to maintain semantic consistency across a wider range of applications. We also intend to integrate self-supervised or semi-supervised learning strategies into MS-Mix to alleviate the reliance on large annotated datasets and enhance its applicability in low-resource settings. Finally, from an efficiency perspective, we will focus on developing lightweight variants of the sample selection and fusion mechanisms to reduce computational costs and support deployment on resource-constrained devices.
6. Acknowledgments
This work was supported by the Science and Technology Innovation Key R&D Program of Chongqing (No. CSTB2024TIAD-STX0023, CSTB2023TIAD-STX0015, and CSTB2023TIAD-STX0031), and in part by the Natural Science Foundation of China (No. 62372427).

### v2

@@ 4.4 Verification Performance of MS-Mix
+ Manifold Mix 78.34±0.44 64.33±0.95 34.94±2.44 78.00±0.78 0.448±0.001 63.90±0.58
+ MultiMix 77.70±1.05 64.80±0.63 36.55±0.99 77.70±0.73 0.447±0.005 64.19±0.27
+ AdvMixUp 78.10±0.54 64.27±0.71 37.61±0.51 78.20±0.47 0.445±0.006 64.55±0.50
+ $\mathcal{P}ow$Mix 78.29±1.03 65.07±0.40 37.74±0.32 78.11±1.06 0.443±0.007 64.80±0.42
+ MS-Mix (ours) 78.51±0.71 65.53±0.43 38.16±0.18 78.70±1.13 0.443±0.003 65.22±0.48
MISA 76.53±0.82 62.96±0.29 36.01±0.64 76.71±0.90 0.452±0.004 63.05±0.44
+ Manifold Mix 76.87±0.69 63.96±0.35 37.67±0.81 77.08±0.69 0.458±0.005 63.89±0.28
+ MultiMix 77.34±0.36 64.26±0.05 37.30±1.18 77.07±1.32 0.439±0.009 63.99±0.21
+ AdvMixUp 77.01±0.80 63.82±0.48 36.47±1.02 77.22±0.52 0.450±0.010 63.63±0.40
+ $\mathcal{P}ow$Mix 77.56±0.39 64.32±0.39 37.21±0.30 77.48±0.33 0.441±0.003 64.14±0.27
+ MS-Mix (ours) 78.19±0.28 64.70±0.59 37.75±0.21 78.14±0.44 0.434±0.004 64.69±0.23
ALMT 77.19±0.88 64.25±0.80 38.68±0.50 76.81±0.86 0.442±0.009 64.23±0.48
+ Manifold Mix 76.88±1.49 63.67±1.00 38.49±1.94 77.08±1.22 0.429±0.004 64.03±0.49
+ MultiMix 77.52±1.30 63.04±0.42 38.62±0.52 78.11±0.09 0.436±0.007 64.32±0.17
+ AdvMixUp 77.84±1.14 64.50±1.30 37.94±0.30 77.88±0.64 0.427±0.003 64.54±0.08
+ $\mathcal{P}ow$Mix 77.61±0.25 64.84±0.77 38.44±0.61 77.87±0.34 0.424±0.004 64.69±0.36
+ MS-Mix (ours) 78.21±0.14 64.52±0.53 38.63±0.41 78.30±0.40 0.426±0.003 64.92±0.36
GLoMo 78.80±0.66 65.66±0.76 37.11±0.86 78.81±0.51 0.427±0.005 65.09±0.51
+ Manifold Mix 78.98±1.16 65.93±0.62 37.78±0.52 78.94±0.67 0.426±0.005 65.41±0.47
+ MultiMix 78.64±0.66 66.42±0.42 38.28±0.41 79.05±0.88 0.428±0.010 65.60±0.28
+ AdvMixUp 78.31±0.88 66.98±0.83 38.30±0.20 77.23±0.36 0.432±0.002 65.20±0.44
+ $\mathcal{P}ow$Mix 79.34±0.65 65.11±1.94 38.33±0.42 79.12±0.18 0.428±0.007 65.47±0.45
+ MS-Mix (ours) 79.33±0.45 67.38±0.98 38.36±0.31 79.23±0.30 0.423±0.001 66.07±0.39
Figure 4: Visualization of loss landscapes comparing Mixup variants and MS-Mix on MOSEI using ALMT.
@@
The consistent outperformance of MS-Mix across diverse backbones and datasets stems from its core design principles, which directly address key limitations in existing mixup-based augmentation methods. Unlike these methods, which often rely on random or offline mixing strategies, MS-Mix incorporates three interconnected mechanisms tailored for MSA: First, the SASS strategy prevents the blending of semantically contradictory samples (e.g., happy and angry) by filtering pairs based on emotional similarity in the latent space. In contrast, the comparative methods [34, 18, 17, 33] mix random sample pairs regardless of emotional coherence, often yielding ambiguous or noisy samples that mislead the training process. By ensuring semantic and label consistency, SASS significantly enhances the quality of augmented data. Second, the SIG mixing module adaptively determines modality-specific mixing ratios based on emotional salience, rather than using random or fixed values as in Manifold Mix [34], MultiMix [18], or leveraging modality importance without emotional context like $\mathcal{P}ow$Mix [17] and AdvMixUp [33]. This allows MS-Mix to dynamically weight each modality’s contribution, leading to more discriminative mixed features and improved robustness of learned cross-modal representations. Third, the SAL serves as a regularization term that aligns predicted sentiment distributions with ground-truth labels, enhancing prediction consistency across modalities and providing additional supervisory signal. This component is unique to MS-Mix and offers a critical advantage over comparative methods, which lack explicit constraints on the alignment of emotion-level distributions. As a result, SAL not only stabilizes training but also strengthens the model’s resistance to overfitting.
@@
Table 4: Results of the ablation experiments using the LMF method on the MOSI database.
Bold: Best performance.
@@
82.02 79.16 82.01 79.27 41.90 36.52 0.879 66.81
SIG → Ground-Truth 79.76 78.29 79.82 78.31 38.27 34.87 0.936 64.89
SIG → Linear Mapping 80.50 78.23 80.79 78.37 39.04 34.70 0.927 65.27
4.5 Performance Visualization
To further demonstrate the effectiveness of MS-Mix, we conduct both t-SNE [48] visualizations (Fig. 3) and loss landscape visualizations (Fig. 4). Specifically, we employ t-SNE to project the distributions of textual features for both the original and mixed features extracted from the MOSEI dataset, using the MISA model. The results reveal that features generated by MS-Mix exhibit significantly clearer decision boundaries and more distinct cluster separation than those produced by $\mathcal{P}$owMix. This enhanced separation indicates that MS-Mix not only preserves the semantic integrity of the original feature space but also yields more discriminative intermediate representations.
Furthermore, we analyze the loss landscape to examine the generalization effect of MS-Mix. The generalization ability of neural networks is closely related to the geometry of their loss landscapes, where convergence to flat minima is generally associated with superior robustness and generalization performance [49]. To visualize this, we plot the 2D loss surface of the ALMT model on MOSEI by perturbing the parameters along two random, mutually orthogonal directions, generated via Gaussian sampling and Gram-Schmidt orthogonalization. As shown in Fig. 4, introducing Manifold Mix and other variants resulted in some landscape smoothing. However, pronounced peaks and high local curvature remained, indicating persistent sensitivity to parameter perturbations. In contrast, MS-Mix yields a notably smoother, flatter landscape, suggesting reduced gradient conflict during training and convergence toward more robust minima.
In summary, by integrating the evidence from label distribution on MOSEI (Fig. 1), feature visualization (Fig. 3), and loss landscapes (Fig. 4), MS-Mix effectively mitigates the semantic confusion and label mismatch issues inherent in other mixup variants. Our approach generates discriminative features with well-matched labels, thereby significantly enhancing the quality of mixed samples and improving overall performance in MSA.
4.6 Ablation and Hyperparameter Analysis
To evaluate the impact of key hyperparameters and individual algorithmic components of MS-Mix, we performed ablation studies on three core components using the LMF [22] model on the MOSI [42] and SIMS [44] datasets. Additionally, hyperparameter sensitivity analyses on the MOSEI dataset [43] were conducted using the MISA [46] and GLoMo [6] models. All experiments were run five times using different random seeds, and the average values were reported as the results.
In our ablation study, we integrated Manifold Mix [34] into the classic LMF [22] backbone as the baseline to evaluate the contribution of each proposed component: SASS, SIG, and SAL. To isolate the effect of each module, we conducted experiments by selectively enabling or disabling individual components during training. Note that SAL relies on SIG’s output and thus cannot be evaluated independently. Furthermore, to justify the rationality of SIG’s complex design, we additionally investigated two simplified variants: degenerating SIG to a linear layer (SIG → Linear Mapping) and using the absolute values of the ground-truth labels as the intensity scores directly (SIG → Ground-Truth). All results were obtained by averaging over five independent runs with different random seeds.
As shown in the table 4, each of the three core components contributes substantially to the overall performance gains over the baseline. Removing any single module results in noticeable degradation, underscoring its individual importance. Furthermore, simplified variants of SIG also lead to clear performance drops. This is because a simple linear mapping is insufficient for extracting the complex, non-linear nature of emotional intensity, and the direct use of a global label overlooks the differences in emotional expression across modalities. In contrast, the attention-based prediction mechanism in our full SIG module effectively captures these latent, modality-specific intensity signals.
@@ 4.6 Ablation and Hyperparameter Analysis
Figure 5: The impact of different parameter values. (a) Accuracy under different similarity thresholds $\delta$. (b) Accuracy under different combinations of $\xi_{1}$ and $\xi_{2}$. (c) Accuracy under different $\alpha$. (d) Accuracy under different attention heads $h$.
@@
To comprehensively analyze the sensitivity of key hyperparameters, we conducted extensive experiments using both the top-performing GLoMo model [6] and the classical MISA architecture [46] on the largest English dataset MOSEI [43] and the Chinese benchmark SIMS [44]. As illustrated in Fig. 5, our evaluation shows that hyperparameters are critical to model performance. Using the MISA model on MOSEI, we first examined the similarity selection threshold $\delta$ (Fig. 5(a)). The results demonstrate that at $\delta=0.2$, the method maintains an optimal balance, incorporating sufficient samples while effectively filtering out emotionally contradictory ones, thus achieving peak performance. Notably, even under suboptimal $\delta$ configurations, MS-Mix consistently surpasses Manifold Mix [34]. We further evaluated the loss weights $\xi_{1}$ and $\xi_{2}$ with the GLoMo model on MOSEI (Fig. 5(b)). The experimental outcomes confirm the rationality of our weight selection, showing that $\xi_{1}=0.7$ and $\xi_{2}=0.5$ yield the best performance. Additionally, we assessed the shape parameter $\alpha$ of the Beta distribution using MISA on MOSEI (Fig. 5(c)), finding $\alpha=2.0$ to be optimal. The influence of the number of attention heads $h$ was also investigated across datasets (Fig. 5(d)): $h=4$ delivers the best results on MOSEI, while $h=6$ is most effective on SIMS. These findings collectively validate the rationale behind our parameter configurations.
4.7 Occlusion Experiment
@@
In the random occlusion experiment, we compared the performance of MS-Mix against baseline methods [34, 18, 17] under both clean and noisy data conditions across MOSI [42], MOSEI [43], and SIMS [44]. Specifically, we randomly masked out 0% to 40% of the input data during training of the LMF [22] model to investigate the effect of mixup-based augmentation on model robustness.
@@
Figure 6: The performance of different mixup-based methods at different occlusion ratios on three datasets.
@@
As shown in Fig. 6, the experimental results indicate that introducing noise affects performance across all datasets. A noteworthy phenomenon emerges: moderate levels of random masking (e.g., 10% or 20%) unexpectedly improve performance rather than degrade it. Beyond these optimal masking ratios, however, model performance deteriorates. Despite these variations, MS-Mix consistently achieves the highest average accuracy across all evaluated conditions, demonstrating its advantage in optimizing the model to handle noisy data.
5 CONCLUSION
@@
In this paper, we propose MS-Mix, a novel, adaptive data augmentation framework designed specifically for MSA. Unlike traditional mixup-based methods, which often rely on random and offline mixing strategies, MS-Mix introduces three key emotion-aware mechanisms to enhance the semantic consistency of augmented samples. First, the SASS module ensures that only emotionally consistent samples are combined, effectively avoiding the mixing of feature pairs with contradictory emotions. Second, the SIG mixing module adaptively determines mixing ratios for each modality based on emotional salience, thereby promoting samples with clearer decision boundaries. Third, the SAL aligns the predicted emotion distributions with ground-truth labels via KL-divergence minimization, serving as an effective regularizer that improves generalization. An empirical evaluation across three benchmark datasets and six diverse backbone architectures shows that MS-Mix achieves state-of-the-art performance.
@@
From an engineering perspective, MS-Mix enables more reliable and data-efficient multimodal emotion understanding, facilitating practical deployment in human-computer interaction, intelligent human-robot systems, and emotion-aware assistive technologies. In future work, we will consider extending MS-Mix to other multimodal tasks, such as mental health monitoring and human-computer interaction, leveraging its ability to maintain semantic consistency in a wider range of applications. Furthermore, we also intend to integrate self-supervised or semi-supervised learning strategies into MS-Mix to reduce reliance on large annotated datasets.
6 Conflicts of Interest Statement
The authors declare no conflicts of interest.